OpenStudy (yeezus):

Let f(x)={cos(pi)(x), x is greater than or equal to -2 {x^2+k, x>2 find a value of k so that this function is continuous

OpenStudy (anonymous):

replace \(x\) by \(-2\) in both expressions set them equal, solve for \(k\)

OpenStudy (anonymous):

You probably mean, x^2 + k, where x < 2. If thats the case, then lets see what happens at x = 2, cos (2pi) = 1 x^2 + k = 4 + k, so for f to be cont. at x = 2, 1 = 4+k Hence, k = - 3.